OpenStudy (anonymous):

add (6x^4+2x^3-9x-7)to(-12x^4-6x^3+21x+32)

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

Combine the numbers that have the same type of x's, like 6x^4 with -12x^4 and that'll give you your x^4. Then you go to x^3 and do those, know what I mean?

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

Let me know what you get so I can check for you :)

OpenStudy (anonymous):

i got -14

OpenStudy (anonymous):

Alright, so another big thing is, keep your terms separate. Keep your x^4 separate from your x^3, they never mix. and don't get rid of those x's either unless they cancel each other out. I'll give you an example. (2x^2-3x+7)+(4x^2+6x-1) You'll start off with the biggest x's 2x^2 and 4x^2 combine them and you get 6x^2 then you go to your next x's, -3x and 6x combine into 3x now combine the last numbers 7 and -1 you get 6 So now you have 6x^2+3x+6

OpenStudy (anonymous):

279

OpenStudy (anonymous):

Here, I'll give you an outline of your problem. Then you solve and put them in order, and does your problem say x=anything? or is what you posted all of it? 6x^4-12x^4 2x^3-6x^3 -9x+21x -7+32

OpenStudy (anonymous):

25

OpenStudy (anonymous):

\[x(-6x^7-4x^5+12x+25)\]

OpenStudy (anonymous):

alright, I'll walk you through it. 6x^4-12x^4 = -6x^4 2x^3-6x^3 = -4x^3 -9x+21x = 12x -7+32 = 25 now you gotta put all your ducks in a row. -6x^4-4x^3+12x+25

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yup :) but I've got an important question. Do you understand what I did?

OpenStudy (anonymous):

yes i did

OpenStudy (anonymous):

i got on my other right thanks for u